# Approximate independence of permutation mixtures

### v2

@@ 5.4 General basis expansions
As mentioned in Section 3, in the Gaussian case where $P_{i}={\mathcal{N}}(\theta_{i},1)$, the methods of moments and cumulants described in Section 2 rely on the expansion $\varphi(x-\theta)/\varphi(x)=1+\sum_{k\geq 1}h_{k}(x)\frac{\theta^{k}}{k!}$, where $\varphi$ is the density of ${\mathcal{N}}(0,1)$ and $h_{k}$ is the $k$th Hermite polynomial. By contrast, the doubly centered expansion is based on the representation $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\phi_{0}(x)=\frac{1}{n}\sum_{i=1}^{n}\varphi(x-\theta_{i})$ denotes the marginal density of each coordinate under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$. This representation possesses the important property that the functions $(x_{1},\dots,x_{n})\mapsto\prod_{i\in S}\Psi(x_{i},\theta_{i})$ are orthogonal with respect to $L^{2}(\mathbb{Q}_{n})$ and $\frac{1}{n}\sum_{i=1}^{n}\Psi(x,\theta_{i})=0$.
@@ Acknowledgments.
We are grateful to Terence Tao for suggesting the use of the saddle point method in the proof of Theorem 3.3, to Cun-Hui Zhang for introducing us the question of the Bayes risk under a permutation prior in Lemma 5.3 and the reference [Zha12], to Yury Polyanskiy for pointing us to the reference [TP23], and to Yunzi Ding and Cheng Mao for discussions about an earlier version of this project. JNW is partially supported by a Sloan Research Fellowship and NSF grant DMS-2210583.
@@ Appendix A Useful identities
We discuss several identities involving permutations and matrix permanents which will be useful in the proofs of several results. Fix an integer $\ell\in[n]$ and $n$ distributions $P_{1},\cdots,P_{n}\in{\mathcal{P}}$. The first quantity appears in (23) of the doubly centered expansion in Section 3, and is formally defined as
@@
with $\overline{P}=\frac{1}{n}\sum_{i=1}^{n}P_{i}$ as usual. The second quantity is the $-\ell$ homogeneous polynomial $S_{\ell}$ in (27) of the matrix permanent approach in Section 4: for the matrix $A\in\mathbb{R}^{n\times n}$ defined in (26) and its eigen-decomposition $A=UDU^{\top}$ with $D=\mathrm{diag}(\lambda_{1},\cdots,\lambda_{n})$, the quantity $S_{\ell}$ is defined as
@@
=(26)${=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{\ell}(nA_{\pi(i),\pi^{\prime}(i)}-1)\right]$
@@
=$\overset{\rm(a)}{=}\frac{n^{\ell}}{\binom{n}{\ell}^{2}\ell!}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{S,S^{\prime}})=\frac{n^{\ell}}{\binom{n}{\ell}^{2}\ell!}\cdot T_{\ell},$
@@
is a polynomial in $t$. An alternative expression of $\mathrm{Perm}(UD_{t}U^{\top})$ is also possible: by the structure of leading eigenvalue and eigenvector in Lemma 4.2, $UD_{t}U^{\top}=U(tD-(t-1)\mathrm{diag}(1,0,\ldots,0))U^{\top}=tA-\frac{t-1}{n}J=t\overline{A}+\frac{J}{n}$ for $J:={\bf 1}{\bf 1}^{\top}$ being the all-one matrix. Consequently,
@@
$=\sum_{S,S^{\prime}\subseteq[n]:|S|=|S^{\prime}|}\mathrm{Perm}\left((t\overline{A})_{S,S^{\prime}}\right)\mathrm{Perm}\left((\frac{J}{n})_{[n]\backslash S,[n]\backslash S^{\prime}}\right)$
@@
Appendix B Proof of Theorem 3.3
Since the complex case (24) has been established in Section 3.3, in this section we are devoted to the proof of (25) in the real case. We split the analysis into several steps.
@@ B.1 The case of binary support
=$\overset{\rm(b)}{=}\frac{1}{\pi r^{\ell}}\int_{-1}^{1}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}\frac{\mathrm{d}t}{\sqrt{1-t^{2}}},$
@@ B.2 Reduction to binary support
For the inductive step, assume that $n\geq\ell+1$ and the induction hypothesis (37) holds for all values smaller than $n$. As the constraint set of (37) is compact, the optimization program (37) must admit a maximizer $x^{\star}$. By the method of Lagrangian multipliers (for maximizing $e_{\ell}(x)$ or $-e_{\ell}(x)$), for any maximizer $x^{\star}$, there must exist $\lambda,\mu\in\mathbb{R}$ such that
$e_{\ell-1}\left(x^{\star}\backslash\left\{x_{i}^{\star}\right\}\right)=\lambda x_{i}^{\star}+\mu,\qquad\text{for all }i\in[n].$ (38)
@@
In this case, the rescaled vector $\sqrt{1-n^{-1}}(x^{\star}\backslash\left\{x_{i}^{\star}\right\})$ satisfies the constraint in (37) with $n$ replaced by $n-1$. By the induction hypothesis,
$|e_{\ell}(x^{\star})|=\left(\frac{n}{n-1}\right)^{\frac{\ell}{2}}\left|e_{\ell}\left(\sqrt{1-n^{-1}}(x^{\star}\backslash\left\{x_{i}^{\star}\right\})\right)\right|\leq\left(\frac{n}{n-1}\right)^{\frac{\ell}{2}}\sqrt{10\binom{n-1}{\ell}}\leq\sqrt{10\binom{n}{\ell}},$
@@
$(x_{2}^{\star}-x_{1}^{\star})e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star}\right\})=\lambda(x_{1}^{\star}-x_{2}^{\star})\Longrightarrow e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star}\right\})=-\lambda.$
Similarly we also have $e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{3}^{\star}\right\})=-\lambda$, and a further subtraction gives
$(x_{3}^{\star}-x_{2}^{\star})e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})=0\Longrightarrow e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})=0.$ (39)
Again, subtracting (39) with $e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{4}^{\star}\right\})=0$ leads to
$(x_{4}^{\star}-x_{3}^{\star})e_{\ell-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0\Longrightarrow e_{\ell-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0.$ (40)
@@
$e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})=x_{4}^{\star}e_{\ell-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})+e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\}),$
@@
$e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0.$ (41)
@@
For a real vector $x\in\mathbb{R}^{n}$ and $10\leq k\leq n-$ 1, if $e_{k}(x)=e_{k+1}(x)=0$, then $e_{\ell}(x)=0$ for all $\ell\geq k$.
As $\ell\geq 4$ and $n\geq\ell+1$, Lemma B.2 applied to (40) and (41) gives $\prod_{i=5}^{n}x_{i}^{\star}=e_{n-4}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},x_{4}^{\star}\right\})=0$, i.e. one of $(x_{5}^{\star},\ldots,x_{n}^{\star})$ must be zero. However, it is assumed at the end of Step I that $x_{i}^{\star}\neq 0$ for all $i\in[n]$, a contradiction! So $|\mathrm{supp}(x^{\star})|\leq 3$, as desired.
@@
$\{x_{1},x_{2},x_{3}\}\cap\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},0\}$ ==$\varnothing$.= ∅ . (44)
@@
$e_{\ell}(x^{\prime})-e_{\ell}(x^{\star})$ $=e_{\ell-1}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})(x_{1}+x_{2}+x_{3}-x_{1}^{\star}-x_{2}^{\star}-x_{3}^{\star})$
$\quad+e_{\ell-2}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})(x_{1}x_{2}+x_{2}x_{3}+x_{1}x_{3}-x_{1}^{\star}x_{2}^{\star}-x_{2}^{\star}x_{3}^{\star}-x_{1}^{\star}x_{3}^{\star})$
$\quad+e_{\ell-3}(x^{\star}\backslash\left\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star}\right\})(x_{1}x_{2}x_{3}-x_{1}^{\star}x_{2}^{\star}x_{3}^{\star})$
@@ B.3 Remaining corner cases
In all these scenarios, the inequality (25) of Theorem 3.3 holds. The proof of Theorem 3.3 is therefore complete.
@@ Appendix C Deferred proofs
C.1 Proof of Corollary 1.3
By Theorem 1.2, it suffices to upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ for the given families ${\mathcal{P}}$. While the computation of $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ is typically straightforward, the evaluation of ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ may require some effort. We present a useful lemma to upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$.
@@
Next we prove Corollary 1.3.
@@
it is clear that ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})=\exp(4\mu^{2})-1$. We show that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})=O(\mu\wedge\mu^{2})$. For $\mu\leq 1$, we simply use ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})=O(\mu^{2})$. For $\mu>1$, we split the class ${\mathcal{P}}$ into ${\mathcal{P}}\subseteq\cup_{m\in\mathbb{Z},|m|\leq\mu+1}{\mathcal{P}}_{m}$, with ${\mathcal{P}}_{m}:=\left\{{\mathcal{N}}(\theta,1):m\leq\theta<m+1\right\}$. Clearly ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}}_{m})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}_{m})=e-1=O(1)$, so Lemma C.1 yields
@@
We only need to prove an additional upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq|\Theta|-1$. In fact, let $\Theta=\{\theta_{1},\ldots,\theta_{m}\}$, Lemma C.1 leads to
@@
so that Equation 4 gives the first upper bound. For the second upper bound, we show that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-2\varepsilon$ for general $\varepsilon>0$, so that (4) gives
@@
Therefore, by Lemma C.1, we have
@@
C.2 Proof of Corollary 1.4
For the first inequality (1), it is clear that
$nI(\vartheta_{1};X_{1})-I(\vartheta;X)$ $=nI(\vartheta_{1};X_{1})-\mathbb{E}_{(\vartheta,X)}\left[\log\frac{\prod_{i=1}^{n}\mathrm{d}P_{\vartheta_{i}}}{\mathrm{d}P_{X}}\right]$
$=nI(\vartheta_{1};X_{1})-\mathbb{E}_{(\vartheta,X)}\left[\log\frac{\prod_{i=1}^{n}\mathrm{d}P_{\vartheta_{i}}}{\mathrm{d}Q_{X}}\right]+\mathbb{E}_{P_{X}}\left[\log\frac{\mathrm{d}P_{X}}{\mathrm{d}Q_{X}}\right]$
$=nI(\vartheta_{1};X_{1})-n\mathbb{E}_{(\vartheta_{1},X_{1})}\left[\log\frac{\mathrm{d}P_{\vartheta_{1}}}{\mathrm{d}Q_{X_{1}}}\right]+{\sf KL}(P_{X}\|Q_{X})$
=$\overset{\rm(a)}{=}{\sf KL}(P_{X}\|Q_{X}),$
where $P_{X},Q_{X}$ are the marginal distributions of the random vector $X$ under the permutation prior and the i.i.d. prior of $\vartheta$, respectively, and (a) uses $P_{X_{1}}=Q_{X_{1}}$. So $nI(\vartheta_{1};X_{1})-I(\vartheta;X)\geq 0$, and
$nI(\vartheta_{1};X_{1})-I(\vartheta;X)$ $\leq\log\left(1+\chi^{2}(P_{X}\|Q_{X})\right)$
$\leq\min\left\{10\sum_{\ell=2}^{n}I_{\chi^{2}}(\vartheta_{1};X_{1})^{\ell},I_{\chi^{2}}(\vartheta_{1};X_{1})(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}}))\right\},$
where we use Theorem 1.2 and replace the channel capacity ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ by a tighter quantity $I_{\chi^{2}}(\vartheta_{1};X_{1})$, which is an intermediate step in the proof of Theorem 1.2.
For the second inequality (8), let $Q_{1},\dots,Q_{m}$ be an $\varepsilon$-cover of $\mathrm{conv}({\mathcal{P}})$ under the KL divergence, with a fixed $\varepsilon>0$ and $m=N_{\mathrm{KL}}(\mathrm{conv}({\mathcal{P}}),\varepsilon)$. By the standard entropic upper bound of the mutual information (cf. [YB99] or [PW24, Theorem 32.4]), one has
$I(\theta;X)$ $\leq\mathbb{E}_{\theta}\left[\min_{i\in[m]}{\sf KL}(P_{X|\theta}\|Q_{i}^{\otimes n})\right]+\log m$
=$\overset{\rm(b)}{=}\mathbb{E}_{\theta}\left[{\sf KL}(P_{X|\theta}\|Q_{X|\theta})+\min_{i\in[m]}\mathbb{E}_{P_{X|\theta}}\left[\log\frac{\mathrm{d}Q_{X|\theta}}{\mathrm{d}Q_{i}^{\otimes n}}\right]\right]+\log m$
=$\overset{\rm(c)}{=}\mathbb{E}_{\theta}\left[{\sf KL}(P_{X|\theta}\|Q_{X|\theta})+\min_{i\in[m]}{\sf KL}(Q_{X|\theta}\|Q_{i}^{\otimes n})\right]+\log m$
where in (b) $Q_{X|\theta}$ denotes the i.i.d. distribution $\left(\frac{1}{n}\sum_{i=1}^{n}P_{\theta_{i}}\right)^{\otimes n}$, (c) uses the i.i.d. structures for both $Q_{X|\theta}$ and $Q_{i}^{\otimes n}$, and notes that the marginal distributions $P_{X_{1}|\theta}$ and $Q_{X_{1}|\theta}$ are the same, and (d) follows from Theorem 1.2 and the definition of the KL covering applied to $\frac{1}{n}\sum_{i=1}^{n}P_{\theta_{i}}\in\mathrm{conv}({\mathcal{P}})$. Taking the infimum over $\varepsilon>0$ then completes the proof of (8).
C.3 Proof of Theorem 1.5
Since $P_{X^{n}}$ is exchangeable, it can be written as a mixture $P_{X^{n}}=\mathbb{E}_{\lambda}\left[\nu_{\mathbb{P}}(\theta_{1,\lambda},\ldots,\theta_{n,\lambda})\right]$, where $\nu_{\mathbb{P}}(\theta_{1},\ldots,\theta_{n})$ denotes the discrete probability measure on $\mathbb{R}^{n}$ induced by sampling without replacement:
$\nu_{\mathbb{P}}(\theta_{1},\ldots,\theta_{n})(x)=\frac{1}{n!}\sum_{\pi\in S_{n}}\prod_{i=1}^{n}{\mathbf{1}_{\left\{{x_{i}=\theta_{\pi(i)}}\right\}}}.$
By joint convexity of the $\chi^{2}$-divergence, it suffices to deal with the scenario $P_{X^{n}}=\nu_{\mathbb{P}}(\theta_{1},\ldots,\theta_{n})$, in which case $P_{Y^{n}}$ is a permutation mixture with components $(P_{1},\ldots,P_{n}):=({\mathsf{K}}(\cdot|\theta_{1}),\ldots,{\mathsf{K}}(\cdot|\theta_{n}))$. The marginal distribution $P_{Y^{k}}$ is then
$\mathrm{d}P_{Y^{k}}(y^{k})=\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\mathrm{d}P_{\pi(i)}(y_{i})\right].$
We will show that the product distribution $Q_{Y^{k}}=\overline{P}^{\otimes k}$ achieves the theorem statement, where as usual $\overline{P}:=\frac{1}{n}\sum_{i=1}^{n}P_{i}$. Specifically, the $\chi^{2}$-divergence is
$\chi^{2}(P_{Y^{k}}\|Q_{Y^{k}})+1$ $=\mathbb{E}_{Q_{Y^{k}}}\left[\left(\frac{\mathrm{d}P_{Y^{k}}}{\mathrm{d}Q_{Y^{k}}}\right)^{2}\right]$
@@
$=\mathbb{E}_{Y_{1},\ldots,Y_{k}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(Y_{i})\right]\right)^{2}\right]$
@@
=$\overset{\rm(a)}{=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\mathbb{E}_{Y_{i}\sim\overline{P}}\left[\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(Y_{i})\frac{\mathrm{d}P_{\pi^{\prime}(i)}}{\mathrm{d}\overline{P}}(Y_{i})\right]\right]$
@@
=$\overset{\rm(b)}{=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[n^{k}\prod_{i=1}^{k}A_{\pi(i),\pi^{\prime}(i)}\right]$
@@
=$\overset{\rm(c)}{=}\frac{n^{k}}{\binom{n}{k}^{2}k!}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\mathrm{Perm}(A_{S,S^{\prime}}),$
@@
where in (a) we introduce an independent copy $\pi^{\prime}\sim\text{\rm Unif}(S_{n})$ of $\pi$, (b) recalls the definition of $A$ in (26), and (c) uses $\mathbb{P}(\pi([k])=S)=\mathbb{P}(\pi^{\prime}([k])=S^{\prime})=\binom{n}{k}^{-1}$ for all subsets $S,S^{\prime}$ of size $k$ and averages over $k!$ bijections between $S$ and $S^{\prime}$. Next we relate $A$ to its centered version $\overline{A}=A-\frac{1}{n}{\bf 1}{\bf 1}^{\top}$. Let $J={\bf 1}{\bf 1}^{\top}$ denote the all-one matrix, then
@@
$=\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\sum_{\begin{subarray}{c}T\subseteq S,T^{\prime}\subseteq S^{\prime}\\ |T|=|T^{\prime}|\end{subarray}}\mathrm{Perm}(\overline{A}_{T,T^{\prime}})\mathrm{Perm}\left(\frac{J_{S\backslash T,S^{\prime}\backslash T^{\prime}}}{n}\right)$

### v3

@@ 6.3 General basis expansions
As mentioned in Section 4, in the Gaussian case where $P_{i}={\mathcal{N}}(\theta_{i},1)$, the methods of moments and cumulants described in Section 3 rely on the expansion $\varphi(x-\theta)/\varphi(x)=1+\sum_{k\geq 1}h_{k}(x)\frac{\theta^{k}}{k!}$, where $\varphi$ is the density of ${\mathcal{N}}(0,1)$ and $h_{k}$ is the $k$th Hermite polynomial. By contrast, the doubly centered expansion is based on the representation $\varphi(x-\theta)/\varphi_{0}(x)=1+\Psi(x,\theta)$, where $\varphi_{0}(x)=\frac{1}{n}\sum_{i=1}^{n}\varphi(x-\theta_{i})$ denotes the marginal density of each coordinate under $\mathbb{P}_{n}$ and $\mathbb{Q}_{n}$. This representation possesses the important property that the functions $(x_{1},\dots,x_{n})\mapsto\prod_{i\in S}\Psi(x_{i},\theta_{i})$ are orthogonal with respect to $L^{2}(\mathbb{Q}_{n})$ and $\frac{1}{n}\sum_{i=1}^{n}\Psi(x,\theta_{i})=0$.
@@ Acknowledgments.
We are grateful to Terence Tao for suggesting the use of the saddle point method for proving Theorem˜4.3, to Cun-Hui Zhang for introducing to us the question of identifying the least favorable prior in Lemma˜2.1 and the reference [Zha12], to Yury Polyanskiy for pointing us to the reference [TP23], to Bero Roos for telling us the references [Bob05, Roo15], and to Yunzi Ding and Cheng Mao for discussions about an earlier version of this project. JNW is partially supported by a Sloan Research Fellowship and NSF grant DMS-2210583.
@@ Appendix A Useful identities
We discuss several identities involving permutations and matrix permanents which will be useful in the proofs of several results. Fix an integer $\ell\in[n]$ and $n$ distributions $P_{1},\cdots,P_{n}\in{\mathcal{P}}$. The first quantity appears in (25) of the doubly centered expansion in Section˜4, and is formally defined as
@@
with $\overline{P}=\frac{1}{n}\sum_{i=1}^{n}P_{i}$ as usual. The second quantity is the degree-$\ell$ homogeneous polynomial $S_{\ell}$ in (29) of the matrix permanent approach in Section˜5: for the matrix $A\in\mathbb{R}^{n\times n}$ defined in (28) and its eigen-decomposition $A=UDU^{\top}$ with $D=\mathrm{diag}(\lambda_{1},\cdots,\lambda_{n})$, the quantity $S_{\ell}$ is defined as
@@
[…]
@@
$\overset{\rm(a)}{=}\frac{n^{\ell}}{\binom{n}{\ell}^{2}\ell!}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=\ell\end{subarray}}\mathrm{Perm}(\overline{A}_{S,S^{\prime}})=\frac{n^{\ell}}{\binom{n}{\ell}^{2}\ell!}\cdot T_{\ell},$
@@
is a polynomial in $t$. An alternative expression of $\mathrm{Perm}(UD_{t}U^{\top})$ is also possible: by the structure of leading eigenvalue and eigenvector in Lemma˜5.2, $UD_{t}U^{\top}=U(tD-(t-1)\mathrm{diag}(1,0,\ldots,0))U^{\top}=tA-\frac{t-1}{n}J=t\overline{A}+\frac{J}{n}$ for $J:={\bf 1}{\bf 1}^{\top}$ being the all-ones matrix. Consequently,
@@
[…]
@@
Appendix B Proof of Theorem˜4.3
Since the complex case (26) has been established in Section˜4.3, in this section we are devoted to the proof of (27) in the real case. We split the analysis into several steps.
@@ B.1 The case of binary support
$\overset{\rm(b)}{=}\frac{1}{\pi r^{\ell}}\int_{-1}^{1}\left(1+\frac{k}{n-k}r^{2}+2r\sqrt{\frac{k}{n-k}}t\right)^{\frac{n-k}{2}}\left(1+\frac{n-k}{k}r^{2}-2r\sqrt{\frac{n-k}{k}}t\right)^{\frac{k}{2}}\frac{\mathrm{d}t}{\sqrt{1-t^{2}}},$
@@ B.2 Reduction to binary support
For the inductive step, assume that $n\geq\ell+1$ and the induction hypothesis (40) holds for all values smaller than $n$. As the constraint set of (40) is compact, the optimization program (40) must admit a maximizer $x^{\star}$. By the method of Lagrangian multipliers (for maximizing $e_{\ell}(x)$ or $e_{\ell}(x)$), for any maximizer $x^{\star}$, there must exist $\lambda,\mu\in\mathbb{R}$ such that
[…]
@@
[…]
@@
[…]
@@
[…]
@@
[…]
@@
For a real vector $x\in\mathbb{R}^{n}$ and $0\leq k\leq n-1$, if $e_{k}(x)=e_{k+1}(x)=0$, then $e_{\ell}(x)=0$ for all $\ell\geq k$.
[…]
@@
$\{x_{1},x_{2},x_{3}\}\cap\{x_{1}^{\star},x_{2}^{\star},x_{3}^{\star},0\}$ $=\varnothing.$ (47)
@@
[…]
@@ B.3 Remaining corner cases
In all these scenarios, the inequality (27) of Theorem˜4.3 holds. The proof of Theorem˜4.3 is therefore complete.
@@ Appendix C Deferred proofs
C.1 Completing the proof of Theorem˜1.2
Since the upper bound (4) of Theorem˜1.2 has been established in Section˜4, we prove the remaining upper bounds (5) and (6) here from Lemma˜5.3. Based on (29) and Lemma˜5.3, we have
Thanks to the spectral gap bound in Lemma˜5.2, we have $\lambda_{i}\in\left[0,1-\frac{1}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}\right]$, so that
$\overset{\rm(a)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\cdot\frac{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})}{\mathsf{\Delta}_{H^{2}}({\mathcal{P}})-1}\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\overset{\rm(b)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\left(1+\log\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\right),$
where (a) follows from Lemma˜5.2, (b) uses $\log x\leq x-1$ for all $x>0$. For the inequality (6), we have
$\overset{\rm(c)}{\leq}{\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\cdot\frac{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1}{{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})}\log({\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1)\overset{\rm(d)}{\leq}({\mathsf{C}}_{\chi^{2}}({\mathcal{P}})+1)\log({\mathsf{D}}_{\chi^{2}}({\mathcal{P}})+1),$
where (c) is due to $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})\leq 1+{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$ and
and (d) uses ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})$. Therefore the upper bounds (5) and (6) are established.
C.2 Proof of Corollary˜1.3
By Theorem˜1.2, it suffices to upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ and $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ for the given families $\mathcal{P}$}. While the computation of $\mathsf{\Delta}_{H^{2}}({\mathcal{P}})$ is typically straightforward, the evaluation of ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$ may require some effort. We present a useful lemma to upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})$.
@@
Next we prove Corollary˜1.3.
@@
it is clear that ${\mathsf{D}}_{\chi^{2}}({\mathcal{P}})=\exp(4\mu^{2})-1$. We show that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})=O(\mu\wedge\mu^{2})$. For $\mu\leq 1$, we simply use ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}})=O(\mu^{2})$. For $\mu>1$, we split the class $\mathcal{P}$} into ${\mathcal{P}}\subseteq\cup_{m\in\mathbb{Z},|m|\leq\mu+1}{\mathcal{P}}_{m}$, with ${\mathcal{P}}_{m}:=\left\{{\mathcal{N}}(\theta,1):m\leq\theta<m+1\right\}$. Clearly ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}}_{m})\leq{\mathsf{D}}_{\chi^{2}}({\mathcal{P}}_{m})=e-1=O(1)$, so Lemma˜C.1 yields
@@
We only need to prove an additional upper bound ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq|\Theta|-1$. In fact, let $\Theta=\{\theta_{1},\ldots,\theta_{m}\}$, Lemma˜C.1 leads to
@@
so that Equation˜4 gives the first upper bound. For the second upper bound, we show that ${\mathsf{C}}_{\chi^{2}}({\mathcal{P}})\leq 1-2\varepsilon$ for general $\varepsilon>0$, so that (4) gives
@@
Therefore, by Lemma˜C.1, we have
@@
C.3 Proof of Theorem˜1.4
Let $\overline{P}:=\frac{1}{n}\sum_{i=1}^{n}P_{i}$. Expanding the likelihood ratio gives
@@
$\chi^{2}(\mathbb{P}_{k,n}\|\mathbb{Q}_{k,n})+1$ $=\mathbb{E}_{X_{1},\ldots,X_{k}\sim\overline{P}}\left[\left(\mathbb{E}_{\pi\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]\right)^{2}\right]$
@@
$\overset{\rm(a)}{=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[\prod_{i=1}^{k}\mathbb{E}_{X_{i}\sim\overline{P}}\left[\frac{\mathrm{d}P_{\pi(i)}}{\mathrm{d}\overline{P}}(X_{i})\frac{\mathrm{d}P_{\pi^{\prime}(i)}}{\mathrm{d}\overline{P}}(X_{i})\right]\right]$
@@
$\overset{\rm(b)}{=}\mathbb{E}_{\pi,\pi^{\prime}\sim\text{\rm Unif}(S_{n})}\left[n^{k}\prod_{i=1}^{k}A_{\pi(i),\pi^{\prime}(i)}\right]$
@@
$\overset{\rm(c)}{=}\frac{n^{k}}{\binom{n}{k}^{2}k!}\sum_{\begin{subarray}{c}S,S^{\prime}\subseteq[n]\\ |S|=|S^{\prime}|=k\end{subarray}}\mathrm{Perm}(A_{S,S^{\prime}}),$
@@
where in (a) we introduce an independent copy $\pi^{\prime}\sim\text{\rm Unif}(S_{n})$ of $\pi$, (b) recalls the definition of $A$ in (28), and (c) uses $\mathbb{P}(\pi([k])=S)=\mathbb{P}(\pi^{\prime}([k])=S^{\prime})=\binom{n}{k}^{-1}$ for all subsets $S,S^{\prime}$ of size $k$ and averages over $k!$ bijections between $S$ and $S^{\prime}$. Next we relate $A$ to its centered version $\overline{A}=A-\frac{1}{n}{\bf 1}{\bf 1}^{\top}$. Let $J={\bf 1}{\bf 1}^{\top}$ denote the all-ones matrix, then
@@
[…]